OpenStudy (anonymous):

Finding zeroes and determining multiplicity help please? For the polynomial function f(x) = (2 - x) * (x + 1) * (x - 3) * (x + 5), find the zeroes. Then determine the multiplicity at each zero and state whether the graph displays the behavior of a touch or a cross at each intercept.

OpenStudy (anonymous):

(A) x = -5, cross; x = -1, cross; x = 2, cross; x = 3, cross (B) x = -5, cross; x = -1, touch; x = 2, touch; x = 3, touch (C) x = -5, cross; x = -1, cross; x = 2, touch; x = 3, touch (D) x = -5, cross; x = -1, cross; x = 2, cross; x = 3, touch

OpenStudy (anonymous):

i'm not sure but i think A is correct if i understand the question correctly

OpenStudy (mathmate):

Hint: 1. make a list of the roots 2. sort them out in ascending order 3. draw a sketch, and compare with the answers.

OpenStudy (mathmate):



OpenStudy (perl):

when it 'touches', the multiplicity is even. when it 'crosses' the multiplicity is odd

OpenStudy (anonymous):

@mathmate im still having trouble

OpenStudy (mathmate):

@JoeJoldin in which step?

OpenStudy (anonymous):

are my roots 2, 1, -3, and 5?

OpenStudy (mathmate):

Close, but not quite. When the polynomial is, say \(f(x)=(x-2)(x+3)\) the roots are such that each factor equal zero, i.e. (x-2)=0, or (x+3)=0, which makes the zeroes x=2 or x-3. Ok so far?

OpenStudy (mathmate):

@JoeJoldin ok so far?

OpenStudy (mathmate):

Can you now list the zeroes?

OpenStudy (anonymous):

@mathmate 5, 1, -2, -3?

OpenStudy (mathmate):

Recall that with the factor (x-2), the zero is obtained by equating to zero (x-2)=0, which, when solved for x, gives x=\( +\)2, the zero for the (x+3) factor is x=-3. So the sign of the zeros that you obtained need to be reversed. You need to sort them in ascending order and proceed to step 3. Since this is a quartic (highest degree is 4, with the leading coefficient negative, the graph looks like:  where the 4 intersections with the x-axis corresponding to the roots. Since there is no repeated roots (such as (x-5)(x-5)=(x-5)^2, it won't be "touching" the x-axis.